# LEMUR: Learned Multi-Vector Retrieval

### v1

@@ Abstract
Multi-vector representations generated by late interaction models, such as ColBERT, enable superior retrieval quality compared to single-vector representations in information retrieval applications. In multi-vector retrieval systems, both queries and documents are encoded using one embedding for each token, and similarity between queries and documents is measured by the MaxSim similarity measure. However, the improved recall of multi-vector retrieval comes at the expense of significantly increased latency. This necessitates designing efficient approximate nearest neighbor search (ANNS) algorithms for multi-vector search. In this work, we introduce Lemur, a simple-yet-efficient framework for multi-vector similarity search. Lemur consists of two consecutive problem reductions: We first formulate multi-vector similarity search as a supervised learning problem that can be solved using a one-hidden-layer neural network. Second, we reduce inference under this model to single-vector similarity search in its latent space, which enables the use of existing single-vector ANNS methods for speeding up retrieval. In addition to performance evaluation on ColBERTv2 embeddings, we evaluate Lemur on embeddings generated by modern multi-vector text models and multi-vector visual document retrieval models. Lemur is an order of magnitude faster than earlier multi-vector similarity search methods.
Machine Learning, ICML
@@ 1 Introduction
Embeddings generated by deep neural networks power modern information retrieval (IR) applications, such as document or passage retrieval, open-ended generation, and question answering. In the single-vector paradigm, both the query and the documents are represented by (single) points in the same vector space. The inner product between a query embedding and a document embedding can be used to measure the document-query similarity in an efficient fashion. The latency of single-vector retrieval can be further improved by leveraging approximate nearest neighbor search (ANNS) libraries, such as Faiss (Douze et al., 2025), ScANN (Guo et al., 2020), and LoRANN (Jääsaari et al., 2024). This enables scaling single-vector retrieval to massive document collections.
@@
Khattab and Zaharia (2020) introduced ColBERT, whose late interaction mechanism enables multi-vector retrieval. Multi-vector models represent both queries and documents as sets of embeddings, generating one embedding per token. In multi-vector retrieval, the similarity between a query and a document is measured by MaxSim similarity
$\mathrm{MaxSim}(X,C)=\sum_{x\in X}\underset{c\in C}{\max}\,\langle x,c\rangle,$
@@
where $X,C\subset\mathbb{R}^{d}$ denote the sets of embeddings of a query and a document, respectively.
@@
Due to their better expressivity enabled by more fine-grained token-level representations, multi-vector models tend to have superior accuracy compared to single-vector models in IR applications (e.g., Khattab et al., 2021; Thakur et al., 2021; Lin et al., 2023). This has motivated the rapid development of novel multi-vector models. After the introduction of ColBERT and ColBERTv2 (Santhanam et al., 2022b), better-performing (Jha et al., 2024; Amini et al., 2025; Clavié, 2025) and more memory-efficient (Takehi et al., 2025) multi-vector models have been introduced. In addition to retrieval from text corpora, multi-vector models such as ColPali (Faysse et al., 2025) have recently achieved state-of-the-art results in visual document retrieval (Günther et al., 2025; Xu et al., 2025; Teiletche et al., 2025).
@@
Figure 1: A schematic overview of the query process (for indexing, see Sec. 3) in the LEMUR framework: The latent representations $\psi(x)$ of the token-level embeddings $x\in X$ are retrieved from the hidden layer of an MLP trained to estimate the MaxSim similarities between a query and each document. The single-vector representation $\Psi(X)$ is obtained by pooling these latent representations. The $k^{\prime}$ most similar documents to $\Psi(X)$ are retrieved using an ANNS index. The final top-$k$ documents are selected by evaluating the exact MaxSim similarities to these $k^{\prime}$ documents.
@@
The improvement in the retrieval accuracy of multi-vector models comes at the expense of increased retrieval latency. This has motivated the development of multi-vector similarity search algorithms and systems, such as PLAID (Santhanam et al., 2022b), DESSERT (Engels et al., 2023), EMVB (Nardini et al., 2024), and IGP (Bian et al., 2025). These methods rely on token-level pruning of documents as the first step of their pipeline: they first retrieve the most similar document tokens from the collection of all the document tokens for each query token, and then limit the search to the documents containing the selected tokens. However, token-level similarity between a query token and a document token is an inaccurate proxy for the MaxSim similarity of the query-document pair containing them (Lee et al., 2023; Jayaram et al., 2024). Thus, a large set of documents has to be reranked to reach accurate results, which leads to prohibitively high latency at high recall levels.
@@
MUVERA (Jayaram et al., 2024) does not rely on token-level pruning. Instead, it reduces multi-vector similarity search to single-vector similarity search by generating a single fixed-dimensional encoding (FDE) per document or query. This single-vector reduction has enabled integrating MUVERA directly into state-of-the-art vector databases such as Weaviate and Qdrant. However, high-dimensional FDEs are required for accurate retrieval (see Fig. 2), which leads to large memory consumption and high latency.
@@
Thus, more efficient multi-vector similarity search methods are required to realize the full potential of multi-vector retrieval. To this end, we introduce a simple-yet-effective framework (see Fig. 1 above) for multi-vector similarity search. The framework consists of two consecutive problem reductions: first to a supervised learning task, and then to single-vector similarity search.
@@
Our first key idea is that the approximation of the MaxSim similarity between a query and the documents can be formulated as a supervised learning task. Specifically, it is a multi-output regression problem where there are as many outputs as there are documents in a corpus. This enables training a neural network directly for this learning task.
@@
Our second key idea is that the MaxSim estimates of the model trained for this task are computed as inner products between two vectors that can be interpreted as single-vector representations of the query and the document. This enables reducing multi-vector similarity search to single-vector similarity search in the latent space of the model, and thus leveraging highly optimized single-vector ANNS libraries to speed up multi-vector retrieval on massive corpora.
@@
We call the proposed framework Lemur (Learned Multi-Vector Retrieval), and publish it as an open source library. In addition to embeddings generated by ColBERTv2, we evaluate the performance of multi-vector similarity search methods on embeddings generated by modern multi-vector text models (Jha et al., 2024; Chaffin, 2025; Clavié, 2025; Amini et al., 2025) and visual document retrieval models (Faysse et al., 2025; Teiletche et al., 2025). Lemur significantly outperforms the state-of-the-art methods on all datasets. The difference is especially pronounced on non-ColBERTv2 embeddings. The code for Lemur is available at https://github.com/ejaasaari/lemur
@@
  • We introduce Lemur, a framework that reduces multi-vector similarity search to single-vector similarity search via a reformulation to a supervised learning task. An instantiation of Lemur consists of a neural network and a single-vector ANNS method.
  • We evaluate the performance of multi-vector similarity search methods on 6 BEIR benchmark datasets (Thakur et al., 2021) embedded using 5 different multi-vector text models, and on the visual retrieval dataset ViDoRe (Loison et al., 2026) embedded using 2 different multi-vector visual document models.
@@ 2.1 Multi-vector retrieval
Multi-vector models represent both the query $\mathcal{X}$ and the $m$ corpus documents $\{\mathcal{C}_{j}\}_{j=1}^{m}$ as sets of vectors by generating a $d$-dimensional contextualized embedding for each query and document token. Denote the query encoder and the document encoder of a multi-vector model by $Q$ and $D$, respectively. Further, denote the multi-vector representation of a query by $X=Q(\mathcal{X})\subset\mathbb{R}^{d}$ , and the multi-vector representations of the documents by $C_{j}=D(\mathcal{C}_{j})\subset\mathbb{R}^{d}$ for each $j=1,\dots,m$. ColBERTv2 and most other multi-vector models truncate the query or augment it by padding it with [MASK] tokens so that all encoded queries have the same number of embeddings, whereas the encoded documents have a variable number of embeddings.
@@ 2.2 Single-vector similarity search
In single-vector similarity search, also called approximate nearest neighbor search (ANNS), both the query and the documents are represented by single embeddings. Denote the single-vector representations of the query and the documents by $x\in\mathbb{R}^{d^{\prime}}$ and $\{c_{j}\}_{j=1}^{m}\subset\mathbb{R}^{d^{\prime}}$, respectively. The task is to approximately identify the $k$ documents that are the most similar to the query, i.e., to approximate
@@
where $c^{(1)},\dots,c^{(m)}$ denote the documents ordered with respect to their similarity with query $x$ in the descending order, and $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a similarity measure. The most common similarity measures used are the inner product, cosine similarity, and (negative) Euclidean distance. When the similarity measure $s$ is the inner product, ANNS is often called maximum inner product search (MIPS) (e.g., Guo et al., 2020; Lu et al., 2023; Zhao et al., 2023).
The quality of the approximation $S\subset[m],|S|=k$ is typically measured by
@@
i.e., as the fraction of the $k$ most similar documents correctly identified, and the efficiency is typically measured by query latency or by queries per second (QPS) (Li et al., 2019; Aumüller et al., 2020; Jääsaari et al., 2025).
@@ 2.3 Multi-vector similarity search
Compared to single-vector similarity measures, computing the MaxSim similarity is much more computationally intensive, since it requires evaluating the inner products between each pair of query embeddings and document embeddings. This makes multi-vector similarity search an even more performance-critical component in multi-vector retrieval. In multi-vector similarity search, the task is to approximate
@@
where $C^{(1)},\dots,C^{(m)}$ denote the documents ordered w.r.t. their similarity with the query $X$ in descending order. This formulation is similar to the single-vector formulation (2) except now both the query $X\subset\mathbb{R}^{d}$ and the documents $C_{j}\subset\mathbb{R}^{d}$ are sets, and the similarity measure $s(\cdot,\cdot)=\mathrm{MaxSim}(\cdot,\cdot)$ operates on sets instead of vectors.
As in the case of single-vector search, we measure the quality of the approximation $S\in[m],|S|=k$, by recall (3) w.r.t. the true MaxSim $k$-nearest neighbors ($k$-$n$), and the efficiency by query latency or by queries per second (QPS).
@@ 3 LEMUR Framework
In Sec. 3.1, we formulate multi-vector similarity search as a supervised learning task, and outline how a neural network can be trained for this task. The key detail is that while the target function $f$ operates on sets, it can be written as a sum of the values of a function $g$ that operates on vectors. Then, in Sec. 3.2, we reduce multi-vector similarity search to single-vector similarity search in the latent space of the model defined in Sec. 3.1. The problem reduction is based on the observation that the MaxSim estimates are computed as inner products between $d^{\prime}$-dimensional vectors at the output layer of the model.
@@ 3.1 Supervised learning formulation
Given a corpus of documents $\{C_{j}\}_{j=1}^{m}$, we consider the task of approximating the MaxSim similarities between a query $X$ and the documents. We use the approximate MaxSim similarities to select $k^{\prime}>k$ candidate documents, and then rerank them by computing the exact MaxSim similarities.
@@
for each $l\in[m]$. While the input of the target function $f$ is a set with variable size, it can be written as
@@
where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ with
$g_{l}(x):=\underset{c\in C_{l}}{\max}\,\langle c,x\rangle$
for each $l\in[m]$. Thus, it suffices to estimate the function $g$ that operates on vectors and represents the contributions of a token-level embedding $x\in X$ to the sum (1).
Estimating $g$ is a multi-output regression problem with $m$ outputs, and an MLP $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ can be directly learned to solve this problem. We use the MSE loss, and train an MLP composed of a feature encoder $\psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$ and a linear output layer (without a bias) so that the function $g$ can be estimated as
@@
where $\Psi(X):=\sum_{x\in X}\psi(x)$ denotes the pooled features of the query.
@@ 3.2 Problem reduction to single-vector ANNS
$f(X)\approx W\Psi(X)=(\langle w_{1},\Psi(X)\rangle,\dots\langle w_{m},\Psi(X)\rangle),$ (6)
where the vectors $w_{j}\in\mathbb{R}^{d^{\prime}}$, $j=1,\dots,m$, are rows of the weight matrix $W$. In other words, each of the $m$ outputs is given by the inner product of the query and a document, both represented in the same $d^{\prime}$-dimensional latent space: the document embeddings $\{w_{j}\}_{j=1}^{m}$ are learned as the weights of the output layer, and the query embedding $\Psi(X)$ is obtained by pooling the feature encoder values $\psi(x)$.
Further, we do not need MaxSim estimates for all $m$ documents but are only interested in finding the $k^{\prime}$ documents with the largest MaxSim estimates. Finding the $k^{\prime}$ largest values in (6) is a just a single-vector MIPS problem in the $d^{\prime}$-dimensional latent space. Hence, given the feature encoder values, MaxSim similarities between the query and the documents can be approximated by a single-vector ANNS algorithm with the weight vectors $\{w_{j}\}_{j=1}^{m}$ as the corpus.

### v2

@@ Abstract
Multi-vector representations generated by late interaction models, such as ColBERT, enable superior retrieval quality compared to single-vector representations in information retrieval applications. In multi-vector retrieval systems, both queries and documents are encoded using one embedding per token, and similarity between queries and documents is measured by the MaxSim similarity measure. However, the improved quality of multi-vector retrieval comes at the expense of significantly increased search latency. In this work, we introduce Lemur, a simple yet efficient framework for multi-vector similarity search. Lemur consists of two consecutive problem reductions: First, we formulate multi-vector similarity search as a supervised learning problem that can be solved using a one-hidden-layer neural network. Second, we reduce inference under this model to single-vector similarity search in its latent space, enabling the use of existing single-vector search indexes to accelerate retrieval. Lemur is an order of magnitude faster than prior multi-vector similarity search methods. Our code is available at https://github.com/ejaasaari/lemur
multi-vector retrieval, late interaction, vector search, approximate nearest neighbor search
@@ 1 Introduction
Embeddings generated by deep neural networks power modern information retrieval (IR) applications, including passage retrieval, document retrieval, and open-domain question answering. In the single-vector paradigm, each query and document is represented by a single vector in a shared vector space. The inner product between the query and document embeddings can be used to measure query-document similarity in an efficient fashion.
The latency of single-vector retrieval can be further reduced by leveraging approximate nearest neighbor search (ANNS) indexes and libraries, such as Faiss (Douze et al., 2026), HNSW (Malkov and Yashunin, 2018), DiskANN (Jayaram Subramanya et al., 2019), ScaNN (Guo et al., 2020), and LoRANN (Jääsaari et al., 2024). This enables scaling single-vector retrieval to massive document collections.
@@
Khattab and Zaharia (2020) introduced ColBERT, whose late interaction mechanism enables multi-vector retrieval. Multi-vector models represent both queries and documents as sets of embeddings, with one embedding per token. In multi-vector retrieval, the similarity between a query and a document is measured by MaxSim similarity
$\mathrm{MaxSim}(X,C)=\sum_{x\in X}\max_{c\in C}\,\langle x,c\rangle,$
@@
where $X,C\subset\mathbb{R}^{d}$ denote the query and document embedding sets, respectively.
@@
Due to their greater expressiveness enabled by fine-grained token-level representations, multi-vector models tend to have superior accuracy compared to single-vector models in IR applications (e.g., Khattab et al., 2021; Thakur et al., 2021; Lin et al., 2023). This has motivated the rapid development of new multi-vector models. After the introduction of ColBERT and ColBERTv2 (Santhanam et al., 2022b), better-performing (Jha et al., 2024; Amini et al., 2025; Clavié, 2025; Chaffin, 2025; Chaffin et al., 2026) and more memory-efficient (Takehi et al., 2025) multi-vector models have been introduced. In addition to retrieval from text corpora, multi-vector models such as ColPali (Faysse et al., 2025) have recently achieved state-of-the-art results in visual document retrieval (Günther et al., 2025; Xu et al., 2025; Teiletche et al., 2025; Moreira et al., 2026).
@@
Figure 1: A schematic overview of the query-time process used by the Lemur framework (for indexing, see Section 3). The latent representations $\psi(x)$ of the token-level embeddings $x\in X$ are obtained from the hidden layer of an MLP trained to estimate the MaxSim similarities between a query and each document. The single-vector representation $\Psi(X)$ is obtained by pooling these latent representations. The $k^{\prime}$ documents most similar to $\Psi(X)$ are retrieved using an ANNS index. The final top-$k$ documents are selected by evaluating the exact MaxSim similarities to these $k^{\prime}$ documents.
@@
The improvement in the retrieval accuracy of multi-vector models comes at the cost of higher retrieval latency. This has motivated the development of multi-vector similarity search algorithms and systems, such as PLAID (Santhanam et al., 2022b), DESSERT (Engels et al., 2023), EMVB (Nardini et al., 2024), and IGP (Bian et al., 2025). These methods rely on token-level pruning of documents as the first step of their pipelines. For each query token, they first retrieve the most similar document tokens from the collection of all document tokens, and then restrict the search to documents containing the selected tokens. However, token-level similarity between a query token and a document token is an inaccurate proxy for the MaxSim similarity of the corresponding query-document pair (Lee et al., 2023; Jayaram et al., 2024). Consequently, a large candidate set must be reranked to obtain accurate results.
@@
MUVERA (Jayaram et al., 2024) does not rely on token-level pruning. Instead, it reduces multi-vector similarity search to single-vector similarity search by generating a single fixed-dimensional encoding (FDE) for each document and query. This single-vector reduction has enabled integrating MUVERA into widely used vector databases, including Weaviate and Qdrant. However, high-dimensional FDEs are required for accurate retrieval (see Figure 2), leading to large memory consumption and high latency.
@@
In this paper, we introduce a simple yet effective framework (see Figure 1) for multi-vector similarity search. The framework is a fast corpus-specific search reduction that consists of two successive problem reductions: first to a supervised learning task and then to single-vector similarity search.
@@
Our first key idea is that approximating the MaxSim similarity between a query and the corpus documents can be formulated as a supervised learning task. More specifically, it is a multi-output regression problem where there are as many outputs as there are documents in the corpus. This enables training a neural network directly for this task.
@@
Our second key idea is that the MaxSim estimates of the model trained for this task are computed as inner products between two vectors that can be interpreted as single-vector representations of the query and the document. This makes it possible to reduce multi-vector similarity search to single-vector similarity search in the latent space of the model and to leverage highly optimized single-vector ANNS libraries to speed up multi-vector retrieval on massive corpora. The design of Lemur enables fast indexing and makes it applicable in both CPU- and GPU-based retrieval systems.
@@
We call the proposed framework Lemur (Learned Multi-Vector Retrieval). In addition to ColBERTv2 embeddings, we evaluate the performance of multi-vector similarity search methods on embeddings generated by modern multi-vector text models (Jha et al., 2024; Chaffin, 2025; Clavié, 2025; Amini et al., 2025) and visual document retrieval models (Faysse et al., 2025; Teiletche et al., 2025). Lemur significantly outperforms the state-of-the-art methods on all datasets. The difference is especially pronounced on embeddings from models other than ColBERTv2.
@@
  • We introduce Lemur, a fast corpus-specific search reduction that reduces multi-vector similarity search over a corpus to single-vector similarity search by reformulating the problem as a supervised learning task.
  • We evaluate the performance of multi-vector similarity search methods on 8 BEIR benchmark datasets (Thakur et al., 2021) embedded using 5 different multi-vector text models, and on the ViDoRe V3 visual retrieval dataset (Loison et al., 2026) embedded using 2 different multi-vector visual document retrieval models.
@@ 2.1 Multi-Vector Retrieval
Multi-vector models represent both the query $\mathcal{X}$ and $m$ corpus documents $\{\mathcal{C}_{j}\}_{j=1}^{m}$ as sets of vectors by generating a $d$-dimensional contextualized embedding for each query and document token. Denote the query encoder and the document encoder of a multi-vector model by $Q$ and $D$, respectively. Further, denote the multi-vector representation of a query by $X=Q(\mathcal{X})\subset\mathbb{R}^{d}$, and the multi-vector representations of the documents by $C_{j}=D(\mathcal{C}_{j})\subset\mathbb{R}^{d}$ for each $j=1,\dots,m$. ColBERTv2 and most other multi-vector models truncate the query or pad it with [MASK] tokens so that all encoded queries have the same number of embeddings, whereas document representations have a variable number of embeddings.
@@ 2.2 Single-Vector Similarity Search
In single-vector similarity search, also called approximate nearest neighbor search (ANNS), both the query and the documents are represented by single embedding vectors. Denote the single-vector representations of the query and the documents by $x\in\mathbb{R}^{d}$ and $\{c_{j}\}_{j=1}^{m}\subset\mathbb{R}^{d}$, respectively. The task is to identify the $k$ documents that are the most similar to the query, i.e., to approximate
@@
where $c^{(1)},\dots,c^{(m)}$ denote the documents ordered with respect to their similarity to the query $x$ in descending order, and $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a similarity measure. When $s$ is the inner product, the corresponding problem is often called maximum inner product search (MIPS) (e.g., Guo et al., 2020; Lu et al., 2023; Zhao et al., 2023).
The quality of the approximation $S\subset[m]$, $|S|=k$, is typically measured by recall:
@@
This is the fraction of the $k$ most similar documents correctly identified. Efficiency is typically measured by query latency or by queries per second (QPS) (Li et al., 2019; Aumüller et al., 2020; Jääsaari et al., 2025).
@@ 2.3 Multi-Vector Similarity Search
Compared to single-vector similarity measures, computing the MaxSim similarity is much more computationally expensive, since it requires evaluating the inner products between every query embedding and every document embedding. This makes multi-vector similarity search an even more performance-critical component in multi-vector retrieval. In multi-vector similarity search, the task is to approximate
@@
where $C^{(1)},\dots,C^{(m)}$ denote the documents ordered with respect to their similarity to the query $X$ in descending order. This formulation is similar to the single-vector formulation (2), except now both the query $X\subset\mathbb{R}^{d}$ and the documents $C_{j}\subset\mathbb{R}^{d}$ are sets, and the similarity measure $s(\cdot,\cdot)=\mathrm{MaxSim}(\cdot,\cdot)$ operates on sets instead of vectors.
As in single-vector search, we measure the quality of the approximation $S\subset[m]$, $|S|=k$, by recall (3) with respect to the true MaxSim $k$-nearest neighbors, and we measure efficiency by query latency or by queries per second (QPS).
@@ 3 LEMUR Framework
In Section 3.1, we formulate multi-vector similarity search as a supervised learning task and outline how an MLP can be trained for this task. In Section 3.2, we then reduce multi-vector similarity search to single-vector similarity search in the learned latent space of the MLP.
@@ 3.1 Supervised Learning Formulation
Given a corpus represented by $\{C_{j}\}_{j=1}^{m}$, we consider the task of approximating the MaxSim similarities between a query $X$ and the documents. We use the approximate MaxSim similarities to select $k^{\prime}>k$ candidate documents and then rerank them by exact MaxSim similarities.
@@
for each $l\in[m]$. While the input of the target function $f$ is a variable-size set, the function can be written as
@@
where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is defined by
$g_{l}(x):=\underset{c\in C_{l}}{\max}\,\langle x,c\rangle$
for each $l\in[m]$. [^1] The function $g_{l}$ is commonly called the support function (Rockafellar, 1997) of $C_{l}$. The support function of a finite set of points is always a convex piecewise-linear function. Thus, it suffices to estimate the function $g$ that operates on vectors and represents the contribution of a token-level embedding $x\in X$ to the sum (1).
Estimating $g$ is a multi-output regression problem with $m$ outputs, and an MLP $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ can be directly learned to solve this problem. We use the mean squared error (MSE) loss and train an MLP composed of a feature encoder $\psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$ and a linear output layer (without a bias) so that the function $g$ can be estimated as
@@
where $\Psi(X):=\sum_{x\in X}\psi(x)$ denotes the pooled query feature vector.
@@ 3.2 Problem Reduction to Single-Vector ANNS
$f(X)\approx(\langle w_{1},\Psi(X)\rangle,\dots,\langle w_{m},\Psi(X)\rangle),$ (6)
where the vectors $w_{j}\in\mathbb{R}^{d^{\prime}}$, $j=1,\dots,m$, are the rows of the weight matrix $W$. In other words, each of the $m$ outputs is given by the inner product of a query embedding and a document embedding, both represented in the same $d^{\prime}$-dimensional latent space: the document embeddings $\{w_{j}\}_{j=1}^{m}$ are learned as the weights of the output layer, and the query embedding $\Psi(X)$ is obtained by pooling the feature encoder values $\psi(x)$.
Further, we do not need MaxSim estimates for all $m$ documents. Instead, we only need to find the $k^{\prime}$ documents with the largest MaxSim estimates. Finding the $k^{\prime}$ largest values in (6) is just a single-vector MIPS problem in the $d^{\prime}$-dimensional latent space. Hence, given the pooled query embedding $\Psi(X)$, the documents with the largest MaxSim estimates can be retrieved using a single-vector ANNS algorithm with the weight vectors $\{w_{j}\}_{j=1}^{m}$ as the corpus.
Algorithm 1 summarizes the query-time pipeline in Lemur.
Using the standard universal approximation theorem for MLPs (Cybenko, 1989; Hornik, 1991), it follows trivially (see Appendix B) that for every $\varepsilon>0$, there exists a hidden dimension $d^{\prime}$ such that $\|f(X)-W\Psi(X)\|_{\infty}\leq|X|\varepsilon$ for any query $X$. Equivalently, for every document $l\in[m]$, $|\mathrm{MaxSim}(X,C_{l})-\langle w_{l},\Psi(X)\rangle|\leq|X|\,\varepsilon$.
Algorithm 1 Lemur Query Process
0: Query embeddings $X$, feature encoder $\psi$, ANNS index built on $\{w_{j}\}_{j=1}^{m}$, document embeddings $\{C_{j}\}_{j=1}^{m}$, number of reranking candidates $k^{\prime}$, output size $k$
0: Top-$k$ documents for query $X$
[…]
3:for all $j\in\mathcal{I}$ do
4:  $s_{j}\leftarrow\mathrm{MaxSim}(X,C_{j})$
5:end for
6:return the $k$ documents in $\mathcal{I}$ with the largest $\{s_{j}\}_{j\in\mathcal{I}}$